# Dense SAE Latents Are Features, Not Bugs

### v1

@@ 4.4 Alphabet Latents
We identify a class of dense latents that selectively boost or suppress large sets of tokens sharing the same initial letter. Unlike prior work that linked latents to the current token’s first letter (chanin2024absorptionstudyingfeaturesplitting, ), these instead relate to the next token’s initial character.
To discover these latents systematically, we examine each latent’s top 100 positive and negative logit contributions by projecting its decoder weights onto the vocabulary space. Then, we collect the corresponding tokens, and select latents where either set contains at least 90% of tokens starting with the same character (excluding the space character “_”). At layer 25, this procedure yields 114 such latents, of which 21 have activation density >0.1, accounting for 20% of all dense latents. These latents span a range of antipodality scores and activation densities, but notably appear as high-density features only at the model’s final layer. We provide some examples from this layer in Table 2.
Interestingly, we observe multiple latents for each letter, varying in specificity: some target a broad set of short tokens sharing only the first letter (e.g., “b” or “c”), while others focus on longer tokens sharing a multi-letter prefix (e.g., “br” or “cu”). We attribute this granularity to feature splitting (bricken2023monosemanticity, ) possibly driven by n-gram frequency, which yields latents with differing activation densities. These latents illustrate how SAEs dedicate dense units to encode output-specific signals related to next-token lexical structure.
@@ 4.5 Meaningful-Word Latents
Figure 7: Meaningful-Word Latents. (Left) AUCs of predicting feature firing, from whether the POS tag is within the specific category. “Meaning word” and “noun/propernoun” are good predictors, while other categories like “verb” are less predictive. (Middle) Example of L2: pair 15089 (blue), 13092 (red) firing pattens on a document, where 15089 fires on “meaning-heavy” words while 13092 fires on proper nouns and functional words (the, in, a). (Right) Example of L3: 7507 firing patterns, where it fires selectively on proper nouns.
The next class of latents that we investigate are those whose firing can be well predicted by the part-of-speech (POS) tag of the token. We create a reduced set of high-level tags from the Brown Corpus (francis1979brown, ) by combining similar tags (e.g., combining plural and singular forms of nouns), [^7] See Table 4 in § E.6 for our full mapping. and capture dense latent activations on 10k sentences ($\approx$ 200k tokens) from the corpus. Then, for each latent, we calculate the AUC-ROC of predicting the binary latent activations given the binary vector of whether a token is within the high-level POS category. Intuitively, this AUC reflects how well the interpretable linguistic category predicts the latent.
@@ 4.6 PCA Latents
Since the top principal components (PCs) are a large fraction of the variance of the activations, one might expect an SAE to learn dense latents that simply reconstruct this subspace. However, we find that this hypothesis is only partly the case: as shown in Figure 20, an antipodal pair of latents consistently reconstruct most of the first PC (cosine similarity $>0.75$), but other latents do not have a large norm percentage in the top PC, even up to the top 5 PC components. The top PC-aligned latents are generally not immediately interpretable and do not fall into any of our classes above. Interestingly, decreasing or increasing the SAE $L_{0}$ and dictionary size does not get rid of PC-aligned latents nor result in significantly more of them (Figure 21).
@@ 4.7 Layer-wise Dynamics
As noted in the taxonomy of dense latents above, and visualized in Figure 2, each class of dense latents is found in specific layer ranges. Dense latents in early layers have more token-dependent activations and track positional information, those in middle layers represent more conceptual directions, and those in the final layers are mostly mechanisms that the model uses to control its output. Inspired by these observations, in this section we further examine layer-wise characteristics of dense latents.
@@
First, we study how the number of dense latents changes across different layers of the model. Figure 8a illustrates the fraction of latents exceeding density thresholds of 0.05, 0.1, 0.2, and 0.3 at each layer. In the early layers (0-4), we observe transient spikes in latents just above the 0.05 and 0.1 thresholds. These latents are largely the part-of-speech related latents in § 4.5. The absence of similar spikes at the 0.2 and 0.3 thresholds suggest that these early fluctuations arise from SAE training variability rather than fundamental differences in the information encoded at different points of the model’s residual stream. Across the middle layers (5–23), the fraction of dense latents is remarkably stable for all thresholds. Finally, the model’s last two layers exhibit an increase in the number of dense latents, indicating a final emergence of dense features prior to unembedding.
@@
We next ask whether the subspace spanned by dense latents remains stable across layers or varies over the model. For each pair of layers, we compute the principal angles between the subspaces defined by latents with density $>0.2$, then take the median angle as a summary statistic: values near 0° indicate largely overlapping subspaces, while values near 90° indicate dissimilarity. Figure 8c visualizes these median angles for every layer pair of Gemma 2 2B. [^8] We find that using a slightly higher density threshold ($>0.2$) makes the subspace similarity pattern more pronounced. The same plot with a lower threshold ($>0.1$) is shown in § A.3, showing the same clustering trend but with reduced overall similarity. Three clusters emerge. Layers 0-4 share a common dense subspace (low angles). This shifts in the middle of the model (layers 10–22), where a new stable subspace persists (mutually low angles). Finally, the last few layers exhibit a pronounced change (large angles relative to earlier layers), consistent with the rise of alphabet and nullspace latents before the unembedding.
@@ 5 Related Work
Transformer models are thought to represent features as linear directions in activation space (mikolov, ; nips2016_a486cd07, ; elhage2021mathematical, ; nanda-etal-2023-emergent, ; pmlr-v235-park24c, ; olah2024lrh, ), with many more features than neurons, leading to superposition (olah2020zoom, ; elhage2022superposition, ). Early work explored sparse dictionary learning to interpret these representations (OLSHAUSEN19973311, ; faruqui-etal-2015-sparse, ; arora-etal-2018-linear, ; zhang2021wordembeddingvisualizationdictionary, ). More recently, sparse autoencoders (SAEs; ng2011sparse, ) have emerged as a scalable and effective implementation of sparse dictionary learning for transformer-based models (yun-etal-2021-transformer, ; bricken2023monosemanticity, ; huben2024sparse, ; rajamanoharan2024improving, ; rajamanoharan2024jumpingaheadimprovingreconstruction, ; kissane2024interpreting, ; bussmann2025learningmultilevelfeaturesmatryoshka, ) that can recover meaningful and causally important features (templeton2024scaling, ; gao2025scaling, ; marks2025sparse, ).
@@
As SAEs have gained traction, recent work has focused on interpreting the meaning of their latent features (chanin2024absorptionstudyingfeaturesplitting, ; leask2025sparse, ). Building on the neuron interpretation methodology in (bills2023language, ), several recent works interpret SAE latents systematically with LLMs (templeton2024scaling, ; paulo2024automaticallyinterpretingmillionsfeatures, ) or automated interpretation strategies (self-explaining, ). A recurring observation across multiple studies are dense latents, which activate on more than 10% or even 50% of tokens (topkvsgated, ; rajamanoharan2024jumpingaheadimprovingreconstruction, ). Whether these latents reflect meaningful internal computations or arise as undesirable artifacts was up until our work an open question (anthropic-interpretable-dense-features, ; removing-dense-latents, ).
@@
A few prior works identify dense LLM representations: two studies (gurnee2024universalneuronsgpt2language, ; chughtai2024understanding, ) identify positional features in LLMs, and another study (stolfo2024confidence, ) finds entropy directions that are always active. Both of these feature types are in our taxonomy of SAE latents.
@@ 6 Discussion, Limitations & Conclusion
Our work shows that dense SAE latents discover intrinsically dense features in the underlying language model representations. This challenges recent efforts that aim to remove dense latents with add-hoc penalties in the SAE loss function (removing-dense-latents, ). Our results motivate future feature-extraction mechanisms that are able to find features that are not necessarily sparse. For example, such techniques might include SAE designs that allocate autoencoder capacity for representing dense subspaces, approaches that optimize circuit sparsity, or techniques like APD (braun2025interpretability, ) that focus on parameter sparsity.
@@
Despite consistently observing the antipodality trend across both TopK and JumpReLU SAEs and across models (Gemma 2 2B and GPT-2 Small), our interpretability analysis primarily focuses on JumpReLU SAEs trained on Gemma-2-2B, using a single dictionary size and sparsity constraint per layer. Future work could broaden analysis to more models, SAE architectures, and SAE sparsities.
@@ Acknowledgments
We would like to express our gratitude to Arthur Conmy, Neel Nanda, and Vilém Zouhar for their valuable feedback and insightful discussions at different point during the development of this project. AS acknowledges the support of armasuisse Science and Technology through a CYD Doctoral Fellowship.
@@ A.1 Antipodal Pairing in Different SAEs
Figure 9, we report antipodality scores (computed as in Eq. 1) for dense latents in three additional SAEs: two TopK SAEs that we trained on the residual streams of GPT-2 (layer 11) and Gemma 2 2B (layer 25), and a JumpReLU SAE from the Gemma Scope suite trained on an earlier layer (16). In all cases, we observe the same trend highlighted in Figure 1c: high-density latents cluster at high antipodality scores, forming near-antipodal pairs that reconstruct specific directions in residual space.
@@ A.2 Dense Latents During Training
In Figure 10, we visualize the number of dense latents (activation frequency $>0.1$) over training steps for each SAE configuration in our ablation experiment described in § 3.1. All curves converge within the first $\sim$100k steps and remain stable throughout training. This early plateau suggests that dense latents are not a product of late-stage optimization noise, but rather emerge early and persist, indicating that they reflect consistent structure in the residual stream rather than transient artifacts.
@@ A.3 Angles Between Residual Stream Subspaces
In Figure 11, we provide further analysis of the evolution of dense latent subspaces across layers. Panel (a) shows the median principal angle between the subspaces spanned by latents with density $>0.1$ at each pair of layers in Gemma 2 2B. These results follow the trend observed in Figure 8c (based on a $>0.2$ cutoff), revealing distinct subspace clusters in the early, middle, and late layers. However, the overall similarity between subspaces is lower here, reflecting the greater variability introduced by including moderately dense latents (density 0.1-0.2).
@@ A.4 Pairwise Similarity Between Latents’ Weights
In Figure 11c, we report for each latent $i$, the maximum-magnitude cosine similarity of its encoder and decoder weights with any other latent $j$. In particular, we show $\operatorname{sim}(\mathbf{W}_{\mathrm{enc}}^{(i)},\mathbf{W}_{\mathrm{enc}}^{(j)})$ and $\operatorname{sim}(\mathbf{W}_{\mathrm{dec}}^{(i)},\mathbf{W}_{\mathrm{dec}}^{(k)})$, where $j=\operatorname*{arg\,max}_{l\neq i}(|\operatorname{sim}(\mathbf{W}_{\mathrm{enc}}^{(i)},\mathbf{W}_{\mathrm{enc}}^{(l)})|)$ and $k=\operatorname*{arg\,max}_{l\neq i}(|\operatorname{sim}(\mathbf{W}_{\mathrm{dec}}^{(i)},\mathbf{W}_{\mathrm{dec}}^{(l)})|)$. We find that the antipodality score $s$ approaches 1 only when both encoder and decoder similarities are close to $-1$.
@@ Appendix B Experimental Details
For the experiment in § 3.1, we trained TopK SAEs [7] on the residual stream activations at layer 25 of Gemma 2 2B using 1 billion tokens from the OpenWebText corpus [17]. Training followed the default configuration of the Sparsify library, [^9] https://github.com/EleutherAI/sparsify and experiment tracking was conducted using Weights & Biases. [^10] https://wandb.ai The ablation experiment on nullspace latents described in Figure 5 was performed on a 10k-token subset of the C4 corpus [19]. Analyses throughout the paper were conducted using the Gemma Scope SAEs [12] with 16k latents trained on the residual stream of Gemma 2 2B. All experiments were implemented in PyTorch [50], with model inspection tools from the TransformerLens library [51]. Data processing used NumPy [52] and Pandas [53], and figures were generated with Plotly [54].

### v2

@@ 4.4 Alphabet Latents
We identify a class of dense latents that selectively boost or suppress large sets of tokens sharing the same initial letter. Unlike prior work that linked latents to the current token’s first letter (chanin2024absorptionstudyingfeaturesplitting), these instead relate to the next token’s initial character.
To discover these latents systematically, we examine each latent’s top 100 positive and negative logit contributions by projecting its decoder weights onto the vocabulary space. Then, we collect the corresponding tokens, and select latents where either set contains at least 90% of tokens starting with the same character (excluding the space character “_”). At layer 25, this procedure yields 114 such latents, of which 21 have activation density >0.1, accounting for 20% of all dense latents. These latents span a range of antipodality scores and activation densities, but notably appear as high-density features only at the model’s final layer. We provide some examples from this layer in Table˜2.
Interestingly, we observe multiple latents for each letter, varying in specificity: some target a broad set of short tokens sharing only the first letter (e.g., “b” or “c”), while others focus on longer tokens sharing a multi‐letter prefix (e.g., “br” or “cu”). We attribute this granularity to feature splitting (bricken2023monosemanticity) possibly driven by n‑gram frequency, which yields latents with differing activation densities. These latents illustrate how SAEs dedicate dense units to encode output-specific signals related to next-token lexical structure.
@@ 4.5 Meaningful-Word Latents
Figure 8: Meaningful-Word Latents. (Left) AUCs of predicting feature firing, from whether the POS tag is within the specific category. “Meaningful word” and “noun/propernoun” are good predictors, while other categories like “verb” are less predictive. (Middle) Example of L2: pair 15089 (blue), 13092 (red) firing patterns on a document, where 15089 fires on “meaning-heavy” words while 13092 fires on proper nouns and functional words (the, in, a). (Right) Example of L3: 7507 firing patterns, where it fires selectively on proper nouns.
The next class of latents that we investigate are those whose firing can be well predicted by the part-of-speech (POS) tag of the token. We create a reduced set of high-level tags from the Brown Corpus (francis1979brown) by combining similar tags (e.g., combining plural and singular forms of nouns), [^7] See Table 3 in § E.6 for our full mapping. and capture dense latent activations on 10k sentences ($\approx$ 200k tokens) from the corpus. Then, for each latent, we calculate the AUC-ROC of predicting the binary latent activations given the binary vector of whether a token is within the high-level POS category. Intuitively, this AUC reflects how well the interpretable linguistic category predicts the latent.
@@ 4.6 PCA Latents
Since the top principal components (PCs) are a large fraction of the variance of the activations, one might expect an SAE to learn dense latents that simply reconstruct this subspace. However, we find that this hypothesis is only partly the case: as shown in Figure˜21, an antipodal pair of latents consistently reconstruct most of the first PC (cosine similarity >0.75>0.75), but other latents do not have a large norm percentage in the top PC, even up to the top 5 PC components. The top PC-aligned latents are generally not immediately interpretable and do not fall into any of our classes above. Interestingly, decreasing or increasing the SAE $L_{0}$ and dictionary size does not eliminate PC-aligned latents nor result in significantly more of them (Figure˜22).
@@ 4.7 Layer-wise Dynamics
As noted in the taxonomy of dense latents above, and visualized in Figure˜3, each class of dense latents is found in specific layer ranges. Dense latents in early layers have more token-dependent activations and track positional information, those in middle layers represent more conceptual directions, and those in the final layers are mostly mechanisms that the model uses to control its output. Inspired by these observations, in this section we further examine layer-wise characteristics of dense latents.
@@
First, we study how the number of dense latents changes across different layers of the model. Figure˜9a illustrates the fraction of latents exceeding density thresholds of 0.05, 0.1, 0.2, and 0.3 at each layer. In the early layers (0-4), we observe transient spikes in latents just above the 0.05 and 0.1 thresholds. These latents are largely the part-of-speech related latents in §˜4.5. The absence of similar spikes at the 0.2 and 0.3 thresholds suggest that these early fluctuations arise from SAE training variability rather than fundamental differences in the information encoded at different points of the model’s residual stream. Across the middle layers (5–23), the fraction of dense latents is remarkably stable for all thresholds. Finally, the model’s last two layers exhibit an increase in the number of dense latents, indicating a final emergence of dense features prior to unembedding.
@@
We next ask whether the subspace spanned by dense latents remains stable across layers or varies over the model. For each pair of layers, we compute the principal angles between the subspaces defined by latents with density >0.2>0.2, then take the median angle as a summary statistic: values near 0° indicate largely overlapping subspaces, while values near 90° indicate dissimilarity. Figure˜9c visualizes these median angles for every layer pair of Gemma 2 2B. [^8] We find that using a slightly higher density threshold ($0.2$) makes the subspace similarity pattern more pronounced. The same plot with a lower threshold ($0.1$) is shown in § A.3, showing the same clustering trend but with reduced overall similarity. Three clusters emerge. Layers 0-4 share a common dense subspace (low angles). This shifts in the middle of the model (layers 10–22), where a new stable subspace persists (mutually low angles). Finally, the last few layers exhibit a pronounced change (large angles relative to earlier layers), consistent with the rise of alphabet and nullspace latents before the unembedding.
@@ 5 Related Work
Transformer models are thought to represent features as linear directions in activation space (mikolov; nips2016_a486cd07; elhage2021mathematical; nanda-etal-2023-emergent; pmlr-v235-park24c; olah2024lrh), with many more features than neurons, leading to superposition (olah2020zoom; elhage2022superposition). Early work explored sparse dictionary learning to interpret these representations (OLSHAUSEN19973311; faruqui-etal-2015-sparse; arora-etal-2018-linear; zhang2021wordembeddingvisualizationdictionary). More recently, sparse autoencoders (SAEs; ng2011sparse) have emerged as a scalable and effective implementation of sparse dictionary learning for transformer-based models (yun-etal-2021-transformer; bricken2023monosemanticity; huben2024sparse; rajamanoharan2024improving; rajamanoharan2024jumpingaheadimprovingreconstruction; kissane2024interpreting; bussmann2025learningmultilevelfeaturesmatryoshka) that can recover meaningful and causally important features (templeton2024scaling; gao2025scaling; marks2025sparse).
@@
As SAEs have gained traction, recent work has focused on interpreting the meaning of their latent features (chanin2024absorptionstudyingfeaturesplitting; leask2025sparse). Building on the neuron interpretation methodology in (bills2023language), several recent works interpret SAE latents systematically. templeton2024scaling propose a rubric-based evaluation method in which a language model (Claude 3 Opus) scores how well a proposed feature description aligns with the contexts on which the latent activates. Similarly, paulo2024automaticallyinterpretingmillionsfeatures propose a pipeline in which natural language interpretations for SAE latents are matched with different contexts and used by an LLM in different tasks that evaluate how good the interpretations are in predicting activating and non-activating contexts. Other recent efforts explore automated interpretation approaches based on self-interpretation strategies (self-explaining). A recurring observation across multiple studies is dense latents, which activate on more than 10% or even 50% of tokens (topkvsgated; rajamanoharan2024jumpingaheadimprovingreconstruction). anthropic-interpretable-dense-features take the 10 most densely activating latents in a cross-layer Transcoder trained on Claude and attempt to manually interpret them, finding plausible interpretations (e.g., “activates on commas,” “activates on non-terminal tokens in multi-token words”) for 6 of the 10 features. In contrast, removing-dense-latents view dense latents as an undesired phenomenon and propose a frequency-based regularizer to discourage their emergence during training. Whether these latents reflect meaningful internal computations or arise as undesirable artifacts was up until our work an open question.
@@
Prior work has also identified dense signals in language model representations more broadly (i.e., components that encode information consistently across many tokens). gurnee2024universalneuronsgpt2language present a taxonomy of universal neurons that appear across GPT-2 models trained with different seeds. Among these, they identify neurons that encode positional information. chughtai2024understanding similarly identify dense positional features in an SAE trained on GPT-2’s layer 0, though they do not explicitly analyze their activation density. Finally, stolfo2024confidence describe neurons that regulate model confidence by tracking entropy and connect them to a component of the residual stream aligned with the quasi-nullspace of the unembedding matrix.
@@ 6 Discussion, Limitations & Conclusion
Our work shows that dense SAE latents discover intrinsically dense features in the underlying language model representations. This challenges recent efforts that aim to remove dense latents with ad-hoc penalties in the SAE loss function (removing-dense-latents). Our results motivate future feature-extraction mechanisms that are able to find features that are not necessarily sparse. For example, such techniques might include SAE designs that allocate autoencoder capacity for representing dense subspaces, approaches that optimize circuit sparsity, or techniques like APD (braun2025interpretability) that focus on parameter sparsity.
@@
Despite consistently observing the antipodality trend across both TopK and JumpReLU SAEs and across models (Gemma 2 2B and GPT-2 Small), our interpretability analysis primarily focuses on JumpReLU SAEs trained on Gemma 2 2B, using a single dictionary size and sparsity constraint per layer. Future work could broaden analysis to more models, SAE architectures, and SAE sparsities.
@@ Acknowledgments
We would like to express our gratitude to Arthur Conmy, Neel Nanda, and Vilém Zouhar for their valuable feedback and insightful discussions at different points during the development of this project. AS acknowledges the support of armasuisse Science and Technology through a CYD Doctoral Fellowship.
Figure 10: Additional Antipodality Plots. Antipodality scores vs. activation density for (a) TopK SAE on GPT-2 (Layer 11), (b) TopK SAE on Gemma 2 2B (Layer 25), and (c) JumpReLU SAE on Gemma 2 2B (Layer 16). Across all configurations, dense latents tend to have high antipodality scores.
@@ A.1 Antipodal Pairing in Different SAEs
Figure˜10, we report antipodality scores (computed as in Eq.˜1) for dense latents in three additional SAEs: two TopK SAEs that we trained on the residual streams of GPT-2 (layer 11) and Gemma 2 2B (layer 25), and a JumpReLU SAE from the Gemma Scope suite trained on an earlier layer (16). In all cases, we observe the same trend highlighted in Figure˜1c: high-density latents cluster at high antipodality scores, forming near-antipodal pairs that reconstruct specific directions in residual space.
@@ A.2 Dense Latents During Training
In Figure˜12, we visualize the number of dense latents (activation frequency >0.1>0.1) over training steps for each SAE configuration in our ablation experiment described in §˜3.1. All curves converge within the first $\sim$100k steps and remain stable throughout training. This early plateau suggests that dense latents are not a product of late-stage optimization noise, but rather emerge early and persist, indicating that they reflect consistent structure in the residual stream rather than transient artifacts.
@@ A.3 Angles Between Residual Stream Subspaces
In Figure˜11, we provide further analysis of the evolution of dense latent subspaces across layers. Panel (a) shows the median principal angle between the subspaces spanned by latents with density >0.1>0.1 at each pair of layers in Gemma 2 2B. These results follow the trend observed in Figure˜9c (based on a >0.2>0.2 cutoff), revealing distinct subspace clusters in the early, middle, and late layers. However, the overall similarity between subspaces is lower here, reflecting the greater variability introduced by including moderately dense latents (density 0.1-0.2).
@@ A.4 Pairwise Similarity Between Latents’ Weights
In Figure˜11c, we report for each latent $i$, the maximum-magnitude cosine similarity of its encoder and decoder weights with any other latent $j$. In particular, we show $\operatorname{sim}(\mathbf{W}_{\mathrm{enc}}^{(i)},\mathbf{W}_{\mathrm{enc}}^{(j)})$ and $\operatorname{sim}(\mathbf{W}_{\mathrm{dec}}^{(i)},\mathbf{W}_{\mathrm{dec}}^{(k)})$, where $j=\operatorname*{arg\,max}_{l\neq i}(|\operatorname{sim}(\mathbf{W}_{\mathrm{enc}}^{(i)},\mathbf{W}_{\mathrm{enc}}^{(l)})|)$ and $k=\operatorname*{arg\,max}_{l\neq i}(|\operatorname{sim}(\mathbf{W}_{\mathrm{dec}}^{(i)},\mathbf{W}_{\mathrm{dec}}^{(l)})|)$. We find that the antipodality score $s$ approaches 1 only when both encoder and decoder similarities are close to 1-1.
@@ A.5 Similarity with SAE Bias
To investigate the relationship between dense latents and the SAEs’ bias terms, we compute the cosine similarity between each SAE decoder vector and the corresponding layer’s decoder bias. Figure 13 shows the absolute cosine similarity for all latents across layers as a function of activation frequency. We observe a small but distinct group of dense latents (upper-right region) that strongly align with the bias.
@@
Figure 14: Dense-subspace ablations on LLaMA-3.2-1B and GPT-2 Small. For each model’s final layer we train a baseline SAE (blue), retrain after ablating the subspace spanned by dense latents (teal), and retrain after ablating an equally sized subspace of non-dense latents (orange). Only removing the dense-latent subspace collapses the high-density tail.
A.6 Additional Dense-latent Subspace Ablations
Figure˜14 replicates the dense-subspace ablation from §˜3.1 on two additional models: LLaMA-3.2-1B [grattafiori2024llama3herdmodels] and GPT-2 Small [radford2019language]. For each model we (i) train a baseline SAE, (ii) retrain after zero-ablating the subspace spanned by dense latents, and (iii) retrain after ablating an equally sized subspace of randomly chosen non-dense latents. All experiments are run at the final residual-stream layer. In both cases, removing the dense-latent subspace collapses the high-density tail—yielding almost no dense latents—whereas ablating a sparse subspace leaves the distribution essentially unchanged. These replications mirror the Gemma 2 2B result and further support that dense latents reflect an intrinsic residual-stream subspace rather than a training artifact.
@@ Appendix B Experimental Details
For the experiment in §˜3.1, we trained TopK SAEs [gao2025scaling] on the residual stream activations at layer 25 of Gemma 2 2B using 1 billion tokens from the OpenWebText corpus [gokaslan2019openweb]. Training followed the default configuration of the Sparsify library, [^9] https://github.com/EleutherAI/sparsify and experiment tracking was conducted using Weights & Biases. [^10] https://wandb.ai The ablation experiment on nullspace latents described in §˜4.3 was performed on a 10k-token subset of the C4 corpus [c4]. Analyses throughout the paper were conducted using the Gemma Scope SAEs [lieberum2024gemmascopeopensparse] with 16k latents trained on the residual stream of Gemma 2 2B. All experiments were implemented in PyTorch [NEURIPS2019_bdbca288], with model inspection tools from the TransformerLens library [nanda2022transformerlens]. Data processing used NumPy [harris2020array] and Pandas [mckinney-proc-scipy-2010], and figures were generated with Plotly [plotly].